In Quincke’s experiment , the sound detected is changed from a maximum to a minimum when the sliding tube is moved through a distance of 2.50 cm. Find the frequency of sound if the speed of sound in air is 340 m s-1.

Grade:11

1 Answers

Kevin Nash
9 years ago
Distance between tow maximum to a minimum is given by, λ/4 = 2.50 cm ⇒ λ = 10 cm = 10^-1 m We know, V = nx ⇒ n = V/λ = 340/10^-1 = 3400 Hz = 3.4 kHz.